OpenStudy (hkmiu):

Can someone explain to me how to solve this? If cot θ = 1.5 find tan θ

OpenStudy (anonymous):

tan = 1/cot

OpenStudy (homeworksucks):

\[\tan(\theta)=\frac{ 1 }{ \cot( \theta) }\]

OpenStudy (homeworksucks):

A bit fancier

OpenStudy (hkmiu):

would it be .6667?